OpenStudy (anonymous):

integral 9x square root x dx [1,25] please

OpenStudy (anonymous):

\[\int\limits_{1}^{25} 9x \sqrt{x} \]

OpenStudy (matt101):

\[\int\limits_{1}^{25}9x \sqrt{x}dx\] \[=\int\limits\limits_{1}^{25}9x^(3/2)dx\] (that's x^3/2) = (18/5)x^5/2|1->25 Solve.

OpenStudy (anonymous):

is that the final answer?

OpenStudy (matt101):

Well you would have to substitute in the limits but that's the integrated function you use to do it :)

OpenStudy (anonymous):

can i have the final answer please because mine was wrong

OpenStudy (matt101):

Sec...I'll calculate it...

OpenStudy (anonymous):

man ur the best

OpenStudy (matt101):

56232/5 = 11246.4

OpenStudy (matt101):

Np :)